4. In testing for the effectiveness of an antacid compound, 5.0 g hydrochloric acid is mixed with magnesium hydroxide.

a..Write a balanced equation for the reaction. /3

b.Determine the number of moles of magnesium hydroxide required to react with the acid /3

c. If there is 24.0 g magnesium hydroxide in the lab, determine if it will be enough to react with acid. /3

5. While practising stoichiometric calculations, a classmate asks you to check his work. Help your classmate by correcting his solution below; make sure to check for correct mathematical format and calculations. Circle the errors and write the corrections beside them.         /3



1
Expert's answer
2021-04-21T07:23:01-0400

a) 2HCl + Mg(OH)2 = MgCl2 + 2H2O


b) n = m /M

M (HCl) = 36.5 g/mol

n (HCl) = 5.0 / 36.5 = 0.14 mol

n (Mg(OH)2) = 1/2 x n (HCl) = 0.14 / 2 = 0.07 mol


c) n = m/M

M (Mg(OH)2) = 58.32 g/mol

n (Mg(OH)2) = 24.0 / 58.32 = 0.4 mol

The available amount of Mg(OH)2 is enough to react with 5.0 g of HCl.